# MentalChat16K: A Benchmark Dataset for Conversational Mental Health Assistance

### v1

@@ 2. Related Work
Benchmark Datasets Benchmark datasets are crucial for advancing NLP research in mental health. Althoff et al. (2016) presented a large-scale, quantitative study on the SNAP dataset (Althoff et al., 2016), a text-message-based counseling conversation dataset containing over 13 million messages. The PsyQA dataset contains Chinese counseling conversations, and Na et al. used CBT prompts with GPT-3.5-turbo-16k to generate CBT-informed responses (Sun et al., 2021; Na, 2024). CounselChat (Bertagnolli, 2020) includes 3.6k questions and answers from online counseling platforms. The HOPE dataset (Malhotra et al., 2022) includes 12.9k annotated utterances from counseling session videos for dialog-act classification, and the MEMO dataset annotates these for mental health counseling summarization (Srivastava et al., 2022). ChatPsychiatrist’s Psych8K dataset (Liu et al., 2023b) comprises data from 260 real counseling recordings. Our MentalChat16K dataset includes face-to-face or video conference conversations, encompassing verbal and non-verbal interactions. These datasets support advancing NLP applications for mental health (Liu et al., 2023b). For a more comprehensive related work, please refer to Appendix A.2.
@@ 3.1.2. Synthetic Data
To enrich our dataset with diverse therapeutic dialogues, we used the OpenAI GPT-3.5 Turbo (OpenAI, 2024b) model to generate 9,775 question-answer pairs with a customized adaptation of the Airoboros self-generation framework  [^2] https://github.com/jondurbin/airoboros. Under the Airoboros framework, we customized a new prompt (see Table 6) to provide clear instructions to generate patient queries using GPT-3.5 Turbo. These queries were then fed back into GPT-3.5 Turbo to generate corresponding responses. These synthetic conversations covered 33 mental health topics, including Relationships, Anxiety, Depression, Intimacy, Family Conflict, etc. The proportion of each topic (Appendix A.3.2) that typically arises in a counseling session according to the CounselChat (Bertagnolli, 2023) platform was specified in the prompt. This method ensures the synthetic conversations authentically mimic the complexity and diversity of therapist-client interactions, thereby exposes our models to a wide spectrum of psychological conditions and therapeutic strategies.
@@ 3.3. Evaluation
To incorporate human evaluation of the model performance, we invited a senior Postdoc and three Master’s students who possessed interdisciplinary expertise in both computer science and medical sciences to compare the responses generated by the base models, fine-tuned models, and baseline models. For each input question, the participants ranked the responses from the following models: (1) the base model, (2) the base model fine-tuned on synthetic data, (3) the base model fine-tuned on interview data, (4) the base model fine-tuned on both synthetic and interview data, and three baseline models: (5) Samantha-1.1, (6) Samantha-1.2, and (7) ChatPsychiatrist. The responses were ranked from 1 to 7, with 1 being the most effective response and 7 being the least effective. For each model, the final ranking results were calculated as the average ranking across 50 input questions randomly sampled from the 200-question evaluation dataset described in Section 3.2. The difference in scoring scales between LLM-based and human evaluators arises from the nature of their distinct evaluation methodologies. LLM scoring provides objective and granular insight into response quality, while human ranking captures intuitive and comparative judgments, which also help minimize bias and simpliy the evaluation tasks. Together, they offer a holistic view of model performance.
@@ 4.5. Results
LLaMA2 * 7.63$\bullet$ 8.01$\bullet$ 8.46$\bullet$ 8.22$\bullet$ 7.53$\bullet$ 7.63$\bullet$ 8.70$\bullet$ 8.26$\bullet$ 7.69$\bullet$ 7.66$\bullet$ 7.34 $\bullet$ 7.63 $\bullet$ 7.46$\bullet$ 7.95$\bullet$ 3.65
@@
Mixtral-8x7B-Instruct-V0.1 * 7.89$\bullet$ 8.06$\bullet$ 8.78$\bullet$ 8.32$\bullet$ 7.78$\bullet$ 7.75$\bullet$ 8.88$\bullet$ 8.31$\bullet$ 7.86$\bullet$ 7.79$\bullet$ 7.53 7.72$\bullet$ 7.79$\bullet$ 8.04$\bullet$ 1.55
@@
Mixtral-8x7B -V0.1 * 7.88$\bullet$ 8.07$\bullet$ 8.77$\bullet$ 8.28$\bullet$ 7.82$\bullet$ 7.70$\bullet$ 8.85$\bullet$ 8.33$\bullet$ 7.93$\bullet$ 7.72$\bullet$ 7.62$\bullet$ 7.72$\bullet$ 7.76$\bullet$ 8.02$\bullet$ 1.90
@@
Vicuna-V1.5 * 7.66$\bullet$ 8.03$\bullet$ 8.54$\bullet$ 8.25$\bullet$ 7.59$\bullet$ 7.62$\bullet$ 8.70$\bullet$ 8.27$\bullet$ 7.70$\bullet$ 7.58$\bullet$ 7.12$\bullet$ 7.58$\bullet$ 7.37$\bullet$ 7.91$\bullet$ 4.00
@@
Zephyr-Alpha * 7.67$\bullet$ 8.05$\bullet$ 8.55$\bullet$ 8.30 7.60$\bullet$ 7.61 8.71$\bullet$ 8.33$\bullet$ 7.73$\bullet$ 7.66 7.27$\bullet$ 7.58 7.38$\bullet$ 7.99$\bullet$ 2.05
@@
ChatPsychiatrist § 6.46 7.54 6.74 7.48 6.45 7.28 7.98 7.68 6.49 6.88 6.68 7.19 5.54 6.40 5.74
Samantha-V1.11 § 6.81 7.90 7.40 8.12 6.77 7.59 8.20 8.16 6.98 7.57 6.66 7.51 6.43 7.58 4.61
Samantha-V1.2 § 6.89 7.96 7.64 8.02 6.77 7.56 8.35 8.10 7.15 7.59 6.75 7.53 6.54 7.55 4.22
@@ 6. Limitations
Finally, the evaluation framework, though robust, relies heavily on automated assessments and selected human evaluations, which might not fully capture the complexities of real-world counseling efficacy. The use of LLMs as evaluators introduces potential biases, and inconsistencies were observed between different evaluators. Advanced evaluation frameworks like G-Eval (Liu et al., 2023a) and Prometheus ((Kim et al., 2023)) were not utilized and could provide more nuanced assessments in future work.
@@ Acknowledgments
We also gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016791.
@@ A.2. Related Work
Several benchmark datasets have been developed to advance natural language processing (NLP) research in mental health, sourced from both text-based and live counseling conversations. From text-based counseling, Althoff et al. (2016) (Althoff et al., 2016) conducted a large-scale analysis of the SNAP counseling conversation dataset, a foundational dataset for understanding language use and patterns in mental health discussions, though access is required. The PsyQA dataset (Sun et al., 2021) contains Chinese counseling conversations crawled from the web, and Na et al. (Na, 2024) used CBT-centric prompts to guide OpenAI’s GPT-3.5-turbo-16k model in providing CBT-informed responses to questions from this dataset. Additionally, (Bertagnolli, 2020) provides 3.6k questions and answers from online counseling platforms, covering a wide range of mental health topics with responses from licensed therapists. From live counseling conversations, the HOPE dataset (Malhotra et al., 2022) includes 12.9k annotated utterances from publicly available counseling session videos on YouTube, specifically for dialog-act classification. The MEMO dataset (Srivastava et al., 2022) further annotates these same utterances for the task of counseling conversation summarization. Additionally, ChatPsychiatrist’s Psych8K dataset (Liu et al., 2023b) comprises training data from 260 real English counseling recordings. The MentalChat16K dataset stands out from the SNAP dataset by being curated from face-to-face or video conference conversations, which encompass both verbal communication and various forms of non-verbal interaction, unlike text-based messaging. These datasets are instrumental in advancing NLP applications aimed at supporting mental health and well-being.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1454-2/25/08doi: 10.1145/XXXXXX.XXXXXX
@@ 2. Related Work
Benchmark Datasets Benchmark datasets are crucial for advancing NLP research in mental health. Althoff et al. (2016) presented a large-scale, quantitative study on the SNAP dataset (Althoff et al., 2016), a text-message-based counseling conversation dataset containing over 13 million messages. The PsyQA dataset contains Chinese counseling conversations, and Na et al. used CBT prompts with GPT-3.5-turbo-16k to generate CBT-informed responses (Sun et al., 2021; Na, 2024). CounselChat (Bertagnolli, 2023) includes 3.6k questions and answers from online counseling platforms. The HOPE dataset (Malhotra et al., 2022) includes 12.9k annotated utterances from counseling session videos for dialog-act classification, and the MEMO dataset annotates these for mental health counseling summarization (Srivastava et al., 2022). ChatPsychiatrist’s Psych8K dataset (Liu et al., 2023b) comprises data from 260 real counseling recordings. Our MentalChat16K dataset includes face-to-face or video conference conversations, encompassing verbal and non-verbal interactions. These datasets support advancing NLP applications for mental health (Liu et al., 2023b). For a more comprehensive related work, please refer to Appendix A.2.
@@ 3.1.2. Synthetic Data
To enrich our dataset with diverse therapeutic dialogues, we used the OpenAI GPT-3.5 Turbo (OpenAI, 2024b) model to generate 9,775 question-answer pairs with a customized adaptation of the Airoboros self-generation framework  [^2] https://github.com/jondurbin/airoboros. Under the Airoboros framework, we customized a new prompt (see Table 6) to provide clear instructions to generate patient queries using GPT-3.5 Turbo. These queries were then fed back into GPT-3.5 Turbo to generate corresponding responses. These synthetic conversations covered 33 mental health topics, including Relationships, Anxiety, Depression, Intimacy, Family Conflict, etc. The proportion of each topic (Appendix A.3.2) that typically arises in a counseling session according to the CounselChat (Bertagnolli, 2023) platform was specified in the prompt. We additionally conducted a random review of 100 synthetic QA pairs to ensure that each question and answer displayed an authentic counseling style and provided guidance suitable for the domain. This combined approach ensures the synthetic conversations authentically mimic the complexity and diversity of therapist-client interactions, thereby exposes our models to a wide spectrum of psychological conditions and therapeutic strategies.
@@ 3.3. Evaluation
To incorporate human evaluation of the model performance, we invited a senior Postdoc and three Master’s students who possessed interdisciplinary expertise in both computer science and medical sciences to compare the responses generated by the base models, fine-tuned models, and baseline models. For each input question, the participants ranked the responses from the following models: (1) the base model, (2) the base model fine-tuned on synthetic data, (3) the base model fine-tuned on interview data, (4) the base model fine-tuned on both synthetic and interview data, and three baseline models: (5) Samantha-1.1, (6) Samantha-1.2, and (7) ChatPsychiatrist. The responses were ranked from 1 to 7, with 1 being the most effective response and 7 being the least effective. For each model, the final ranking results were calculated as the average ranking across 50 input questions randomly sampled using a fixed seed of 42 to ensure reproducibility from the 200-question evaluation dataset described in Section 3.2. The difference in scoring scales between LLM-based and human evaluators arises from the nature of their distinct evaluation methodologies. LLM scoring provides objective and granular insight into response quality, while human ranking captures intuitive and comparative judgments, which also help minimize bias and simpliy the evaluation tasks. Together, they offer a holistic view of model performance.
@@ 4.5. Results
Discussion on Average Scores The average scores across all 7 metrics evaluated by both GPT and Gemini (14 scores in total for each row) demonstrate the superiority of synthetic data fine-tuning. Among the 7 base models evaluated, models fine-tuned exclusively on synthetic data (*) achieve the highest average scores in 6 cases, with scores ranging from 7.85 to 8.04 compared to 4.02-7.99 for other variants. This consistent advantage observed in diverse architectures, including LLaMA2, Mistral and Mixtral families, indicates that synthetic counseling conversations provide highly effective training data for mental health support capabilities.
@@
Model (7B) Active Listening $\uparrow$ Empathy & Validation $\uparrow$ Safety & Trustworthiness $\uparrow$ Open-mindedness & Non-judgment $\uparrow$ Clarity & Encouragement $\uparrow$ Boundaries & Ethical $\uparrow$ Holistic Approach $\uparrow$ Average Score $\uparrow$ Human Rank $\downarrow$
@@
LLaMA2 * 7.63$\bullet$ 8.01$\bullet$ 8.46$\bullet$ 8.22$\bullet$ 7.53$\bullet$ 7.63$\bullet$ 8.70$\bullet$ 8.26$\bullet$ 7.69$\bullet$ 7.66$\bullet$ 7.34 $\bullet$ 7.63 $\bullet$ 7.46$\bullet$ 7.95$\bullet$ 7.87 3.65
@@
Mixtral-8x7B-Instruct-V0.1 * 7.89$\bullet$ 8.06$\bullet$ 8.78$\bullet$ 8.32$\bullet$ 7.78$\bullet$ 7.75$\bullet$ 8.88$\bullet$ 8.31$\bullet$ 7.86$\bullet$ 7.79$\bullet$ 7.53 7.72$\bullet$ 7.79$\bullet$ 8.04$\bullet$ 8.04 1.55
@@
Mixtral-8x7B -V0.1 * 7.88$\bullet$ 8.07$\bullet$ 8.77$\bullet$ 8.28$\bullet$ 7.82$\bullet$ 7.70$\bullet$ 8.85$\bullet$ 8.33$\bullet$ 7.93$\bullet$ 7.72$\bullet$ 7.62$\bullet$ 7.72$\bullet$ 7.76$\bullet$ 8.02$\bullet$ 8.03 1.90
@@
Vicuna-V1.5 * 7.66$\bullet$ 8.03$\bullet$ 8.54$\bullet$ 8.25$\bullet$ 7.59$\bullet$ 7.62$\bullet$ 8.70$\bullet$ 8.27$\bullet$ 7.70$\bullet$ 7.58$\bullet$ 7.12$\bullet$ 7.58$\bullet$ 7.37$\bullet$ 7.91$\bullet$ 7.85 4.00
@@
Zephyr-Alpha * 7.67$\bullet$ 8.05$\bullet$ 8.55$\bullet$ 8.30 7.60$\bullet$ 7.61 8.71$\bullet$ 8.33$\bullet$ 7.73$\bullet$ 7.66 7.27$\bullet$ 7.58 7.38$\bullet$ 7.99$\bullet$ 7.89 2.05
@@
ChatPsychiatrist § 6.46 7.54 6.74 7.48 6.45 7.28 7.98 7.68 6.49 6.88 6.68 7.19 5.54 6.40 6.91 5.74
Samantha-V1.11 § 6.81 7.90 7.40 8.12 6.77 7.59 8.20 8.16 6.98 7.57 6.66 7.51 6.43 7.58 7.39 4.61
Samantha-V1.2 § 6.89 7.96 7.64 8.02 6.77 7.56 8.35 8.10 7.15 7.59 6.75 7.53 6.54 7.55 7.46 4.22
@@ 6. Limitations
Finally, the evaluation framework, though robust, relies heavily on automated assessments and selected human evaluations, which might not fully capture the complexities of real-world counseling efficacy. The use of LLMs as evaluators introduces potential biases, and inconsistencies were observed between different evaluators. Advanced evaluation frameworks like G-Eval (Liu et al., 2023a) and Prometheus (Kim et al., 2023) were not utilized and could provide more nuanced assessments in future work.
@@
Finally, because our pipeline breaks interviews into isolated QA pairs, MentalChat16K does not preserve multi-turn conversational flow (e.g., follow-up questions, clarifications). In real counseling, exchanges often span multiple turns. We plan to release a multi-turn subset in a future version to support dialog-flow research.
@@ Acknowledgments
We also gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017811.
@@ A.2. Related Work
Several benchmark datasets have been developed to advance natural language processing (NLP) research in mental health, sourced from both text-based and live counseling conversations. From text-based counseling, Althoff et al. (2016) (Althoff et al., 2016) conducted a large-scale analysis of the SNAP counseling conversation dataset, a foundational dataset for understanding language use and patterns in mental health discussions, though access is required. The PsyQA dataset (Sun et al., 2021) contains Chinese counseling conversations crawled from the web, and Na et al. (Na, 2024) used CBT-centric prompts to guide OpenAI’s GPT-3.5-turbo-16k model in providing CBT-informed responses to questions from this dataset. Additionally, (Bertagnolli, 2023) provides 3.6k questions and answers from online counseling platforms, covering a wide range of mental health topics with responses from licensed therapists. From live counseling conversations, the HOPE dataset (Malhotra et al., 2022) includes 12.9k annotated utterances from publicly available counseling session videos on YouTube, specifically for dialog-act classification. The MEMO dataset (Srivastava et al., 2022) further annotates these same utterances for the task of counseling conversation summarization. Additionally, ChatPsychiatrist’s Psych8K dataset (Liu et al., 2023b) comprises training data from 260 real English counseling recordings. The MentalChat16K dataset stands out from the SNAP dataset by being curated from face-to-face or video conference conversations, which encompass both verbal communication and various forms of non-verbal interaction, unlike text-based messaging. These datasets are instrumental in advancing NLP applications aimed at supporting mental health and well-being.